The air quality standard for carbon monoxide is 9 ppm. If we stepped into an air source that contains 9 ppm CO, and took one deep breath, how many molecules of CO would we inhale? [Our total lung capacity is about 1 liter, and it is estimaed that one liter of air contains about 2*10^22 molecules of air]
1
Expert's answer
2019-09-25T05:44:53-0400

During the inhalation, totally 2×1022 molecules of air enter the lungs. This air contains 9 ppm CO molecules. As 1 ppm = 1/1,000,000 = 10-6, the inhaled air contains 2×1022 × 9×10-6 = 18×1016 molecules of carbon monoxide.


Answer: 18×1016 molecules of CO.